OpenStudy (anonymous):

Find the area of a sector of a circle whose diameter is 16 cm and whose central angle is 63 degrees. A. 36.2 cm² B. 35.2 cm² C. 8.8 cm² D. 8.36 cm²

OpenStudy (accessdenied):

Do you know how to find the area of the entire circle?

OpenStudy (accessdenied):

I like to think of sectors as a sort of percentage of the whole circle, like a pie chart or such. The 63 degree central angle is part of 360 whole degrees, so the area of the sector is 63/360 of the total area.

OpenStudy (anonymous):

0.175?

OpenStudy (accessdenied):

That would be multiplied onto the total area of a circle, which is given by \( A = \pi r^2 \)

OpenStudy (accessdenied):



OpenStudy (anonymous):

how do I find the radius ? @AccessDenied

OpenStudy (accessdenied):

You would use the diameter. Diameter is just twice the radius, right? diameter = 16 radius = 1/2 diameter = 1/2 (16)

OpenStudy (anonymous):

Area= 3.14*16^2

OpenStudy (accessdenied):

Radius is 1/2 (16) = 8, so area uses the 8: 3.14 * 8^2.

OpenStudy (anonymous):

3.14*8^2= 200.96 so what do I do next?

OpenStudy (accessdenied):

This is where I mean to use that fraction 63/360. 63 is the part of the circle we are interested in. 360 is the total circle. The fraction of the two is basically the percentage of the circle made up by our sector, so 63/360 or 0.175 or 17.5% of the circle is our sector. So 17.5% of the area is the area of our sector.

OpenStudy (accessdenied):

\( \text{Area of Sector} = \dfrac{63}{360} \times \text{Area of Circle} = 17.5\% \text{ of Area of Circle} \)

OpenStudy (anonymous):

ï don't have that as a answer choice.

OpenStudy (accessdenied):

What did you get? o.o 63/360 * 3.14*8^2

OpenStudy (anonymous):

my bad I miss calculated its B. thanks so much!

OpenStudy (accessdenied):

Yup, glad to help! :)